# Statistical Rounding Error Analysis for the Computation of Random Vectors and Matrices

### v2

@@ IV-B Rounding Error Analysis for LU Factorization
  • For the upper triangular matrix ${\bf U}=\left(u_{kj}\right)$, if $j=k$, the expectation and variance of the rounding errors $\Delta u_{kk}$ are given by
    $\mathbb{E}\left(\Delta u_{kk}\right)$ $=0,$ (25)
    $\mathbb{V}\left(\Delta u_{kk}\right)$ $\approx\left(m^{2}-4\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]-3\left(k-1\right)$
    $+3\sum_{i=1}^{k-1}{\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}}$
    If $j\neq k$, the expectation and variance of the rounding errors $\Delta u_{kj}$ are given by
    $\mathbb{E}\left(\Delta u_{kj}\right)$ $=0,$ (27)
    $\mathbb{V}\left(\Delta u_{kj}\right)$ $\approx\left(m-2\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+\sum_{i=1}^{k-1}{\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}}$
    $-2\frac{\left(1+\sigma^{2}\right)\left[\left(1+\sigma^{2}\right)^{k-2}-1\right]}{\sigma^{2}}+k-3,$ (28)
    where $k\leq j\leq n$, $\sigma_{\epsilon_{i}}^{2}=\left(m-i-1\right)\mathbb{V}\left(\Delta l_{ki}\right),\sigma_{\eta_{i}}^{2}=\frac{\mathbb{V}\left(\Delta u_{ij}\right)}{m-i+1},1\leq i\leq k-1,1\leq k\leq n$.
  • For the lower triangular matrix ${\bf L}=\left(l_{ik}\right)$, the expectation and variance of the rounding errors $\Delta l_{ik}$ are given by
    $\mathbb{E}\left(\Delta l_{ik}\right)$ $=0,$ (29)
    $\mathbb{V}\left(\Delta l_{ik}\right)$ $\approx\frac{(m-6)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}$
    $+\frac{\left(1+\sigma_{\eta_{k}}^{2}\right)\sum_{j=1}^{k-1}{\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}}-k+1}{\left(m-k-1\right)\left(m-k-3\right)}$
    where $k+1\leq i\leq n$, $\sigma_{\eta_{k}}^{2}=\frac{\mathbb{V}\left(\Delta u_{kk}\right)}{\left(m-k+1\right)\left(m-k+3\right)},1\leq k\leq n$.
@@
V Numerical experiments
In this section, we conduct a series of numerical experiments to validate the accuracy of our derivations presented in Sec. III and IV.
@@ V-A Simulation Setup
The experiments are performed using MATLAB R2023b. While most computations are conducted in single precision, Sec. V-B1 employs $\mathtt{fp16}$ and $\mathtt{bfloat16}$ arithmetic. MATLAB $\mathtt{single.m}$ function simulates single precision, and the rounding function $\mathtt{chop.m}$ introduced in [37] is used for simulating $\mathtt{fp16}$ and $\mathtt{bfloat16}$ arithmetic. Exact results for inner products and other matrix computations are obtained in double precision.
@@ V-B1 Inner Products
Different input random vectors.
(a) Random uniform $\mathcal{U}\left(0,1\right)$ vectors.
(b) Random uniform $\mathcal{U}\left(-1,1\right)$ vectors.
(c) Random Gaussian $\mathcal{N}\left(0,1\right)$ vectors.
(d) Random Gaussian $\mathcal{N}\left(1,1\right)$ vectors.
We consider the case where the entries of the input vectors have different distributions. Specifically, uniform $\mathcal{U}\left(0,1\right)$ and $\mathcal{U}\left(-1,1\right)$ distributions, as well as Gaussian $\mathcal{N}\left(0,1\right)$ and $\mathcal{N}\left(1,1\right)$ distributions. As shown in Fig. 1, the comparison between simulated variance and analytical variance, i.e., (6), of the rounding error for the computation in single precision is illustrated. We can find that the analytical and simulated curves are very tight in different distributions, confirming our derived results’ correctness. Moreover, the variance of the rounding error grows exponentially with the input dimension since the rounding errors accumulate along the vector dimension. Note that inputs with zero mean exhibit lower variances than those with nonzero mean.
@@
Then, for the worst-case bounds and using [7, Theorem 3.1, Eq. (3.6)], we have
@@
Next, based on the probabilistic bound in [8, Theorem 3.2], we obtain
@@
Finally, we utilize [9, Theorem 3.2 & 3.3], yielding
@@
Lower precision arithmetic. Now we repeat the experiment conducted in Sec. V-B1 using Gaussian $\mathcal{N}\left(0,1\right)$ distribution, but this time employing precision lower than single precision to compute the inner product $s={\bf x}^{T}{\bf y}$. The simulation results using $\mathtt{fp16}$ and $\mathtt{bfloat16}$ arithmetic are depicted in Fig. 3a and Fig. 3b, respectively. Notably, our analytical results remain highly consistent with the simulated curves even at lower precision, affirming the accuracy of our derivation.
@@ V-C1 Triangular Systems
$\mathbb{V}\left(\Delta x_{3}\right)$ $\approx\frac{\left(1+\sigma^{2}\right)^{3}+\sum_{j=1}^{2}{\mathbb{V}\left(x_{j}\right)\left(1+\sigma_{\psi_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{i-j+2}}}{m-4}-\mathbb{V}\left(x_{3}\right)$
$=\frac{\left(1+\sigma^{2}\right)^{3}+\frac{\left(1+\sigma^{2}\right)^{5}}{m-2}+\frac{\left(m-1\right)\left(1+\sigma_{\psi_{2}}^{2}\right)\left(1+\sigma^{2}\right)^{3}}{\left(m-2\right)\left(m-3\right)}}{m-4}-\mathbb{V}\left(x_{3}\right),$ (38)
@@ VI Conclusions
We have performed the statistical rounding error analysis for random vectors and matrices computation and provided approximate closed-form expressions for the expectation and variance of the rounding errors under a probabilistic model of the relative error. The general analysis has covered inner products, matrix-vector products, and matrix-matrix products. For specific Wishart matrices, we have presented an example of ZF detection and the corresponding LS problem and derived the approximate closed-form expressions for the expectation and variance of the rounding errors for the solution of the triangular system and LU factorization. The numerical experiments have confirmed the accuracy of the analytical expressions and demonstrated that they are generally at least two orders of magnitude tighter than alternative worst-case bounds.
@@ Appendix A Proof of Theorem 1
$\hat{s}_{3}$ $=\boldsymbol{fl}\left(\hat{s}_{2}+\boldsymbol{fl}\left(x_{3}y_{3}\right)\right)=\left(\hat{s}_{1}+x_{3}y_{3}\left(1+\delta_{4}\right)\right)\left(1+\delta_{5}\right)$
@@
Further, we can derive the variance of $\Delta s$ in (43) as follows:
$\mathbb{V}\left(\Delta s\right)$ =$=\left\{\mathbb{V}\left(x_{1}y_{1}\right)+\left[\mathbb{E}\left(x_{1}y_{1}\right)\right]^{2}\right\}\mathbb{V}\left(\left[\prod_{i}^{\left(n\right)}{\left(1+\delta_{i}\right)}-1\right]\right)$
$+\sum_{k=2}^{n}{\left\{\mathbb{V}\left(x_{k}y_{k}\right)+\left[\mathbb{E}\left(x_{k}y_{k}\right)\right]^{2}\right\}\mathbb{V}\left(\left[\prod_{i}^{\left(n-k+2\right)}{\left(1+\delta_{i}\right)}-1\right]\right)}$
@@
$=\tau\left[\left(1+\sigma^{2}\right)^{n}-1\right]+\tau\sum_{k=2}^{n}{\left[\left(1+\sigma^{2}\right)^{n-k+2}-1\right]}$
$+2\sum_{k=2}^{n}{\mathbb{E}\left(x_{1}y_{1}x_{k}y_{k}\left[\prod_{i}^{\left(n-k+1\right)}{\left(1+\delta_{i}^{2}\right)}-1\right]\right)}$
@@
$=\tau\left\{\left[\left(1+\sigma^{2}\right)^{n}-1\right]+\sum_{k=2}^{n}{\left[\left(1+\sigma^{2}\right)^{n-k+2}-1\right]}\right\}$
$+2\mu_{x}^{2}\mu_{y}^{2}\sum_{i=1}^{n-1}{\left[\frac{\left(1+\sigma^{2}\right)\left[\left(1+\sigma^{2}\right)^{n-i}-1\right]}{\sigma^{2}}-n+i\right]}$
@@
$+2\mu_{x}^{2}\mu_{y}^{2}\left[\frac{\left(1+\sigma^{2}\right)^{2}\left[\left(1+\sigma^{2}\right)^{n-1}-1\right]}{\sigma^{4}}-\frac{\left(n-1\right)\left(1+\sigma^{2}\right)}{\sigma^{2}}-\frac{n\left(n-1\right)}{2}\right].$ (43)
where
@@
Then, substituting (44) into (7), neglecting the high-order term of $\sigma^{2}$ and preserving the high-order terms of $n$, we can obtain (8). And the proof ends.
@@ Appendix B Proof of Corollary 1
Using [8, Corollary 3.1], if $\left|x_{i}\right|,\left|y_{i}\right|$ has bounds with means $\mu_{\left|{x}\right|}$ and $\mu_{\left|{y}\right|}$, we can obtain
@@ Appendix C Proof of Theorem 4
$=\frac{b_{i}}{t_{ii}}\left[\prod_{k}^{\left(i\right)}{\left(1+\delta_{k}\right)}-1\right]-\sum_{j=1}^{i-1}{\frac{t_{ij}x_{j}}{t_{ii}}\left[\left(1+\sigma_{\psi_{j}}^{2}\right)\prod_{k}^{\left(i-j+2\right)}{\left(1+\delta_{k}\right)}-1\right]}$
$=z_{i}\left[\prod_{k}^{\left(i\right)}{\left(1+\delta_{k}\right)}-1\right]-\sum_{j=1}^{i-1}{c_{j}x_{j}}\left[\left(1+\sigma_{\psi_{j}}^{2}\right)\prod_{k}^{\left(i-j+2\right)}{\left(1+\delta_{k}\right)}-1\right].$ (60)
@@
Similarly, the variance of the rounding error $\Delta x_{i}$ is given by
$\mathbb{V}\left(\Delta x_{i}\right)$ $=\mathbb{V}\left(z_{i}\right)\mathbb{V}\left(\left[\prod_{k}^{\left(i\right)}{\left(1+\delta_{k}\right)}-1\right]\right)$
$+\sum_{j=1}^{i-1}{\mathbb{V}\left(c_{j}\right)\mathbb{V}\left(x_{j}\right)\mathbb{V}\left(\left[\left(1+\sigma_{\psi_{j}}^{2}\right)\prod_{k}^{\left(i-j+2\right)}{\left(1+\delta_{k}\right)}-1\right]\right)}$
$=\frac{\left(1+\sigma^{2}\right)^{i}+\sum_{j=1}^{i-1}{\mathbb{V}\left(x_{j}\right)\left(1+\sigma_{\psi_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{i-j+2}}}{m-i-1}-\mathbb{V}\left(x_{i}\right),$ (61)
where $\sigma_{\psi_{j}}^{2}$ and $\mathbb{V}\left(x_{j}\right)$ are given in (59) and (56), respectively. And the proof is done.
@@ Appendix D Proof of Theorem 5
$\hat{u}_{kj}$ $=a_{kj}\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-\sum_{i=1}^{k-1}{l_{ki}u_{ij}\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}}.$
@@
$-\sum_{i=1}^{k-1}{q_{i}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]}.$ (67)
Note that the distributions of $a_{kj}$ and $q_{i}$ depend on whether $j=k$ or $j\neq k$. Therefore, if $j=k$, according to Lemma 2, Lemma 3, Lemma 7 and Lemma 8, the expectation of $\Delta u_{kk}$ can be expressed as
@@
and the variance of $\Delta u_{kk}$ is derived as follows:
$\mathbb{V}\left(\Delta u_{kk}\right)$ $=\mathbb{V}\left(a_{kk}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)$
$+\mathbb{V}\left(\sum_{i=1}^{k-1}{q_{i}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]}\right)$
$-2\sum_{i=1}^{k-1}{\mathbb{C}\left(a_{kk}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right],q_{i}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)}$
@@
$-2\sum_{i=1}^{k-1}{\mathbb{E}\left(a_{kk}q_{i}\left[\prod_{r}^{\left(k-i\right)}{\left(1+\delta_{r}^{2}\right)}-1\right]\right)}$
$=\left(m^{2}+2m\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+3\sum_{i=1}^{k-1}{\left[\left(1+\sigma_{\epsilon_{1}}^{2}\right)\left(1+\sigma_{\eta_{1}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}-1\right]}$
$-2\left(m+2\right)\sum_{i=1}^{k-1}{\left[\left(1+\sigma^{2}\right)^{k-i}-1\right]}$
@@
and the expectation of $\Delta u_{kj}$ is determined as follows:
$\mathbb{V}\left(\Delta u_{kj}\right)$ $=\mathbb{V}\left(a_{kj}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)+\mathbb{V}\left(\sum_{i=1}^{k-1}{q_{i}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]}\right)$
$-2\sum_{i=1}^{k-1}{\mathbb{C}\left(a_{kj}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right],q_{i}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)}$
$\approx m\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+\sum_{i=1}^{k-1}{\left[\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}-1\right]}$
$-2\sum_{i=1}^{k-1}{\left[\left(1+\sigma^{2}\right)^{k-i}-1\right]}$
$=\left(m-2\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+\sum_{i=1}^{k-1}{\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}}$
$-2\frac{\left(1+\sigma^{2}\right)\left[\left(1+\sigma^{2}\right)^{k-2}-1\right]}{\sigma^{2}}+k-3.$ (69)
Then, considering the rounding error of $\bf U$, the computed $\hat{l}_{ik},1\leq i\leq k-1$ can be expressed as

### v3

@@ IV-B Rounding Error Analysis for LU Factorization
  • For the upper triangular matrix ${\bf U}=\left(u_{kj}\right)$, if $j=k$, the expectation and variance of the rounding errors $\Delta u_{kk}$ are given in (25) and (26) at the bottom of the page, respectively.
    If $j\neq k$, the expectation and variance of the rounding errors $\Delta u_{kj}$ are given in (27) and (28) at the bottom of the page, where $k\leq j\leq n$, $\sigma_{\epsilon_{i}}^{2}=\left(m-i-1\right)\mathbb{V}\left(\Delta l_{ki}\right),\sigma_{\eta_{i}}^{2}=\frac{\mathbb{V}\left(\Delta u_{ij}\right)}{m-i+1},1\leq i\leq k-1,1\leq k\leq n$.
  • For the lower triangular matrix ${\bf L}=\left(l_{ik}\right)$, the expectation and variance of the rounding errors $\Delta l_{ik}$ are given in (29) and (30) at the bottom of the page, where $k+1\leq i\leq n$, $\sigma_{\eta_{k}}^{2}=\frac{\mathbb{V}\left(\Delta u_{kk}\right)}{\left(m-k+1\right)\left(m-k+3\right)},1\leq k\leq n$.
@@
V Simulation Results and Discussion
(a) Random uniform $\mathcal{U}\left(0,1\right)$ vectors.
(b) Random uniform $\mathcal{U}\left(-1,1\right)$ vectors.
(c) Random Gaussian $\mathcal{N}\left(0,1\right)$ vectors.
(d) Random Gaussian $\mathcal{N}\left(1,1\right)$ vectors.
In this section, we conduct a series of simulation experiments to validate the accuracy of our derivations presented in Section III and IV.
@@ V-A Simulation Setup
The experiments are performed using MATLAB R2023b. While most computations are conducted in single precision, Section V-B1 employs $\mathtt{fp16}$ and $\mathtt{bfloat16}$ arithmetic. MATLAB $\mathtt{single.m}$ function simulates single precision, and the rounding function $\mathtt{chop.m}$ introduced in [41] is used for simulating $\mathtt{fp16}$ and $\mathtt{bfloat16}$ arithmetic. Exact results for inner products and other matrix computations are obtained in double precision.
@@ V-B1 Inner Products
Different input random vectors. We consider the case where the entries of the input vectors have different distributions. Specifically, uniform $\mathcal{U}\left(0,1\right)$ and $\mathcal{U}\left(-1,1\right)$ distributions, as well as Gaussian $\mathcal{N}\left(0,1\right)$ and $\mathcal{N}\left(1,1\right)$ distributions. As shown in Fig. 1, the comparison between simulated variance and analytical variance, i.e., (6), of the rounding error for the computation in single precision is illustrated. We can find that the analytical and simulated curves are very tight in different distributions, confirming our derived results’ correctness. Moreover, the variance of the rounding error grows exponentially with the input dimension since the rounding errors accumulate along the vector dimension. Note that inputs with zero mean exhibit lower variances than those with nonzero mean.
@@
Then, for the worst-case bounds and using [11, Theorem 3.1, Eq. (3.6)], we have
@@
Next, based on the probabilistic bound in [12, Theorem 3.2], we obtain
@@
Finally, we utilize [13, Theorem 3.2 & 3.3], yielding
@@
Lower precision arithmetic. Now we repeat the experiment conducted in Section V-B1 using Gaussian $\mathcal{N}\left(0,1\right)$ distribution, but this time employing precision lower than single precision to compute the inner product $s={\bf x}^{T}{\bf y}$. The simulation results using $\mathtt{fp16}$ and $\mathtt{bfloat16}$ arithmetic are depicted in Fig. 3a and Fig. 3b, respectively. Notably, our analytical results remain highly consistent with the simulated curves even at lower precision, affirming the accuracy of our derivation.
@@ V-C1 Triangular Systems
$\mathbb{V}\left(\Delta x_{3}\right)$ $\approx\frac{\left(1+\sigma^{2}\right)^{3}+\sum_{j=1}^{2}{\mathbb{V}\left(x_{j}\right)\left(1+\sigma_{\psi_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{i-j+2}}}{m-4}$
$-\mathbb{V}\left(x_{3}\right)$
$=\frac{\left(1+\sigma^{2}\right)^{3}+\frac{\left(1+\sigma^{2}\right)^{5}}{m-2}+\frac{\left(m-1\right)\left(1+\sigma_{\psi_{2}}^{2}\right)\left(1+\sigma^{2}\right)^{3}}{\left(m-2\right)\left(m-3\right)}}{m-4}$
$-\mathbb{V}\left(x_{3}\right),$ (38)
@@ VI Conclusions
We have performed the statistical rounding error analysis for random matrix computations and provided approximate closed-form expressions for the expectation and variance of the rounding errors under a probabilistic model of the relative error. The general analysis has covered inner products, matrix-vector products, and matrix-matrix products. For specific Wishart matrices, we have presented an example of ZF detection and the corresponding LS problem and derived the approximate closed-form expressions for the expectation and variance of the rounding errors for the solution of the triangular system and LU factorization. The numerical experiments have confirmed the accuracy of the analytical expressions and demonstrated that they are generally at least two orders of magnitude tighter than alternative worst-case bounds.
@@ Appendix A Proof of Theorem 1
$\hat{s}_{3}$ $=\boldsymbol{fl}\left(\hat{s}_{2}+\boldsymbol{fl}\left(x_{3}y_{3}\right)\right)=\left(\hat{s}_{2}+x_{3}y_{3}\left(1+\delta_{4}\right)\right)\left(1+\delta_{5}\right)$
@@
Further, we can derive the variance of $\Delta s$ in (43) at the bottom of the next page, where
$\mathbb{V}\left(\Delta s\right)$ =$=\left\{\mathbb{V}\left(x_{1}y_{1}\right)+\left[\mathbb{E}\left(x_{1}y_{1}\right)\right]^{2}\right\}\mathbb{V}\left(\left[\prod_{i}^{\left(n\right)}{\left(1+\delta_{i}\right)}-1\right]\right)+\sum_{k=2}^{n}{\left\{\mathbb{V}\left(x_{k}y_{k}\right)+\left[\mathbb{E}\left(x_{k}y_{k}\right)\right]^{2}\right\}\mathbb{V}\left(\left[\prod_{i}^{\left(n-k+2\right)}{\left(1+\delta_{i}\right)}-1\right]\right)}$ (43)
@@
$=\tau\left[\left(1+\sigma^{2}\right)^{n}-1\right]+\tau\sum_{k=2}^{n}{\left[\left(1+\sigma^{2}\right)^{n-k+2}-1\right]}+2\sum_{k=2}^{n}{\mathbb{E}\left(x_{1}y_{1}x_{k}y_{k}\left[\prod_{i}^{\left(n-k+1\right)}{\left(1+\delta_{i}^{2}\right)}-1\right]\right)}$
@@
$=\tau\left\{\left[\left(1+\sigma^{2}\right)^{n}-1\right]+\sum_{k=2}^{n}{\left[\left(1+\sigma^{2}\right)^{n-k+2}-1\right]}\right\}+2\mu_{x}^{2}\mu_{y}^{2}\sum_{i=1}^{n-1}{\left[\frac{\left(1+\sigma^{2}\right)\left[\left(1+\sigma^{2}\right)^{n-i}-1\right]}{\sigma^{2}}-n+i\right]}$
@@
$+2\mu_{x}^{2}\mu_{y}^{2}\left[\frac{\left(1+\sigma^{2}\right)^{2}\left[\left(1+\sigma^{2}\right)^{n-1}-1\right]}{\sigma^{4}}-\frac{\left(n-1\right)\left(1+\sigma^{2}\right)}{\sigma^{2}}-\frac{n\left(n-1\right)}{2}\right].$
@@
Then, substituting (44) in (43), neglecting the high-order term of $\sigma^{2}$ and preserving the high-order terms of $n$, we can obtain (8). And the proof ends.
@@ Appendix B Proof of Corollary 1
Using [12, Corollary 3.1], if $\left|x_{i}\right|,\left|y_{i}\right|$ has bounds with means $\mu_{\left|{x}\right|}$ and $\mu_{\left|{y}\right|}$, we can obtain
@@ Appendix C Proof of Theorem 4
$=\frac{b_{i}}{t_{ii}}\left[\prod_{k}^{\left(i\right)}{\left(1+\delta_{k}\right)}-1\right]$
$-\sum_{j=1}^{i-1}{\frac{t_{ij}x_{j}}{t_{ii}}\left[\left(1+\sigma_{\psi_{j}}^{2}\right)\prod_{k}^{\left(i-j+2\right)}{\left(1+\delta_{k}\right)}-1\right]}$
$=z_{i}\left[\prod_{k}^{\left(i\right)}{\left(1+\delta_{k}\right)}-1\right]$
$-\sum_{j=1}^{i-1}{c_{j}x_{j}}\left[\left(1+\sigma_{\psi_{j}}^{2}\right)\prod_{k}^{\left(i-j+2\right)}{\left(1+\delta_{k}\right)}-1\right].$ (60)
$\mathbb{V}\left(\Delta x_{i}\right)$ $=\mathbb{V}\left(z_{i}\right)\mathbb{V}\left(\left[\prod_{k}^{\left(i\right)}{\left(1+\delta_{k}\right)}-1\right]\right)+\sum_{j=1}^{i-1}{\mathbb{V}\left(c_{j}\right)\mathbb{V}\left(x_{j}\right)\mathbb{V}\left(\left[\left(1+\sigma_{\psi_{j}}^{2}\right)\prod_{k}^{\left(i-j+2\right)}{\left(1+\delta_{k}\right)}-1\right]\right)}$ (61)
$=\frac{\left(1+\sigma^{2}\right)^{i}+\sum_{j=1}^{i-1}{\mathbb{V}\left(x_{j}\right)\left(1+\sigma_{\psi_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{i-j+2}}}{m-i-1}-\mathbb{V}\left(x_{i}\right).$
@@
Similarly, the variance of the rounding error $\Delta x_{i}$ is given in (61) at the bottom of the page, where $\sigma_{\psi_{j}}^{2}$ and $\mathbb{V}\left(x_{j}\right)$ are given in (59) and (56), respectively. And the proof is done.
@@ Appendix D Proof of Theorem 5
$\hat{u}_{kj}$ $=a_{kj}\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}$
$-\sum_{i=1}^{k-1}{l_{ki}u_{ij}\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}}.$
@@
$-\sum_{i=1}^{k-1}{q_{kij}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]}.$ (67)
Note that the distributions of $a_{kj}$ and $q_{kij}$ depend on whether $j=k$ or $j\neq k$. Therefore, if $j=k$, according to Lemma 1, Lemma 2, Lemma 6 and Lemma 7, the expectation of $\Delta u_{kk}$ can be expressed as
@@
and the variance of $\Delta u_{kk}$ is derived in (68) at the bottom of the page.
$\mathbb{V}\left(\Delta u_{kk}\right)$ $=\mathbb{V}\left(a_{kk}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)+\mathbb{V}\left(\sum_{i=1}^{k-1}{q_{kij}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]}\right)$ (68)
$-2\sum_{i=1}^{k-1}{\mathbb{C}\left(a_{kk}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right],q_{kij}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)}$
@@
$-2\sum_{i=1}^{k-1}{\mathbb{E}\left(a_{kk}q_{kij}\left[\prod_{r}^{\left(k-i\right)}{\left(1+\delta_{r}^{2}\right)}-1\right]\right)}$
$=\left(m^{2}+2m\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+3\sum_{i=1}^{k-1}{\left[\left(1+\sigma_{\epsilon_{1}}^{2}\right)\left(1+\sigma_{\eta_{1}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}-1\right]}-2\left(m+2\right)\sum_{i=1}^{k-1}{\left[\left(1+\sigma^{2}\right)^{k-i}-1\right]}$
@@
$-2\left(m+2\right)\left[\frac{\left(1+\sigma^{2}\right)\left[\left(1+\sigma^{2}\right)^{k-2}-1\right]}{\sigma^{2}}-k+2\right].$
@@
and the expectation of $\Delta u_{kj}$ is determined in (69) at the top the next page.
$\mathbb{V}\left(\Delta u_{kj}\right)$ $=\mathbb{V}\left(a_{kj}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)+\mathbb{V}\left(\sum_{i=1}^{k-1}{q_{kij}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]}\right)$ (69)
$-2\sum_{i=1}^{k-1}{\mathbb{C}\left(a_{kj}\left[\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-1\right],q_{kij}\left[\left(1+\epsilon_{i}\right)\left(1+\eta_{i}\right)\prod_{r}^{\left(k-i+1\right)}{\left(1+\delta_{r}\right)}-1\right]\right)}$
$\approx m\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+\sum_{i=1}^{k-1}{\left[\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}-1\right]}-2\sum_{i=1}^{k-1}{\left[\left(1+\sigma^{2}\right)^{k-i}-1\right]}$
$=\left(m-2\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+\sum_{i=1}^{k-1}{\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}}-2\frac{\left(1+\sigma^{2}\right)\left[\left(1+\sigma^{2}\right)^{k-2}-1\right]}{\sigma^{2}}+k-3.$
Then, considering the rounding error of $\bf U$, the computed $\hat{l}_{ik},1\leq i\leq k-1$ can be expressed in (70) at the top of the next page, where we use (63), and define $\eta_{k}=\frac{\hat{u}_{kk}-u_{kk}}{u_{kk}}$ with the mean zero and variance $\sigma_{\eta_{k}}^{2}=\frac{\mathbb{V}\left(\Delta u_{kk}\right)}{\mathbb{V}\left(u_{kk}\right)+\left[\mathbb{E}\left(u_{kk}\right)\right]^{2}}$ at $\left(b\right)$, and $\left(c\right)$ follows the Taylor expansion.
$\Delta l_{ik}$ $\approx o_{ik}\left[\left(1-\eta_{k}\right)\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)-1}\right]$
$-\sum_{j=1}^{k-1}{p_{ijk}\left[\left(1-\eta_{k}\right)\left(1+\epsilon_{j}\right)\left(1+\eta_{j}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}-1\right]}.$